The Grade 11 students decided to have a movie fund-raising activity. It costs 225 pesos per ticket for a student and 500 pesos per ticket for adults. Oddly enough, they lost their tally, but they were able to count the total number of tickets sold (74) and a revenue of 28750. How many student tickets were sold? How many adult tickets were sold?
Answer by richwmiller(17219)

s+a=74
225*s+500*a=28750
We know the total items. 74
x+y=74
We know that the total cost for the two types is 28750
225*x+500*y= 28750
x=74-y
We substitute for x
225*(74-y)+500*y= 28750
We multiply out
16650-225*y+500*y= 28750
We combine like terms.
275*y=28750-16650
275*y=12100
275*y=12100
Isolate y
y=12100/275
y=44 items at $ 500 cost $22000
Calculate x
x=74-44
x=30 items at $ 225 cost $6750
We check
225*30+500*44=28750
6750+22000=28750
28750=28750
Since 28750 = 28750 is TRUE, all is well